OpenStudy (anonymous):

Help with 3 algebra questions?(: Please?(: I'll give a metal(:

OpenStudy (anonymous):

. Which shows 58^2 - 42^2 being evaluated using the difference of squares method? (3 points) 58^2 - 42^2 = (3364 + 1764)(3364 - 1764) = 8,204,800 58^2 - 42^2 = 3025 - 2025 = 1,600 582 - 422 = (58 + 42)(58 - 42) = (100)(16) = 1,600 58^2 - 42^2 = (58 - 42)^2 = 162 = 256 13. Four polynomial functions are shown below: f(x) = x^3 + 3x^2 - 4x - 12 g(x) = x^3 - 3x - 2 p(x) = x^2 + 3x + 2 t(x) = x^2 + x - 2 Which of the following functions has -3, -2, and 2 as its zeros? (4 points) f(x) g(x) p(x) t(x) 14. Which of the following best describes the graph of f(x) = x2 + 7x + 10? (4 points) Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (anonymous):

wait. first one is A

OpenStudy (anonymous):

@pitamar @sleepyjess @sleepyhead314 @jim_thompson5910 @studier97170 @studygeek15

OpenStudy (sleepyjess):

For the second one, we need to put in the zeros as (x + ...) and multiply all 3. Right now we have x = -3, x = -2 and x = 2

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

can you also help with the last one? Just a guide?

OpenStudy (sleepyjess):

Not really sure on the last one, that would be easiest if you used a graphing program to graph it

OpenStudy (sleepyjess):

So what do we need to multiply for the second?

OpenStudy (anonymous):

ok... Thank you(: would the second one be g(x)?

OpenStudy (sleepyjess):

The options didn't show up

OpenStudy (anonymous):

oh dang. let me get them up again

OpenStudy (anonymous):

f(x) = x3 + 3x2 - 4x - 12 g(x) = x3 - 3x - 2 p(x) = x2 + 3x + 2 t(x) = x2 + x - 2

OpenStudy (sleepyjess):

How did you get that?

OpenStudy (anonymous):

through the numbers you showed me. with the x+...

OpenStudy (sleepyjess):

What did you use though? (x + what) for each one?

OpenStudy (anonymous):

just forget I said anything. I did it wrong :/

OpenStudy (sleepyjess):

It's ok :) That's what I'm here for

OpenStudy (anonymous):

sleepyjess to the rescue(:

OpenStudy (anonymous):

14 is C

OpenStudy (anonymous):

13 is f(x)

OpenStudy (anonymous):

12 is C

OpenStudy (sleepyjess):

agree with 13 and 14

OpenStudy (josedavid):

what grade is this becuae if thi sis 9 i could help

OpenStudy (josedavid):

yeap this is 9 i have the same questions im doing it right now

OpenStudy (josedavid):

i think 12 is d

OpenStudy (anonymous):

I already finished algebra @josedavid